OpenStudy (anonymous):

solve for y: (y+3)^(3/2)-8=19

OpenStudy (anonymous):

y = 6.

OpenStudy (anonymous):

\[19 = (y + 3)^{\frac{3}{2}} \]\[27 = (y + 3)^{\frac{3}{2}}\]\[27^{\frac{2}{3}} = y + 3\]\[9 = y + 3\]\[6 = y\]